OpenStudy (anonymous):

The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD?

OpenStudy (skullpatrol):

Do you know the formula for perpendicular slopes?

OpenStudy (skullpatrol):

$$\huge m_1*m_2=-1$$

OpenStudy (anonymous):

It says like 1/2 and 2 and -1/2 and -2 as the answer

OpenStudy (skullpatrol):

What is the slope of line CD?

OpenStudy (anonymous):

The equation of line CD is (y−3) = − 2 (x − 4). It says that

OpenStudy (skullpatrol):

So from that tell me what its slope is please.

OpenStudy (anonymous):

Idk how to do this

OpenStudy (skullpatrol):

This equation (y−3) = − 2 (x − 4) is in what is called the point-slope form: $$y-y_0=m(x-x_0)$$ Have you seen this form before?

OpenStudy (anonymous):

No

OpenStudy (skullpatrol):

m is the slope of this line, please tell what m is equal to.

OpenStudy (anonymous):

-2?

OpenStudy (skullpatrol):